A laser pulse with wavelength 535nm contains 4.40mJ of energy.

How many photons are in the laser pulse?
Answer #1

wavelength = wavespeed/frequency
freq = c/wavelength =3E8/535E-9 = 5.60E14 Hz
photons have energy E = hf = 6.626E-34 Js * 5.60E14 Hz = 3.71 E-19 J
so if it contains 4.40 mJ of energy, the pulse has:
4.40E-3/3.71E-19 = 1.18E16 photons